### v1

@@ Abstract
In a constituent quark model, a hidden charmed tetraquark is assumed consisting of a $cq$ diquark and an $\bar{c}\bar{q}$ antidiquark or vice versa. The Semay-Silvestre-Brac potentials are employed to calculate the masses of $cq$ (q=u, d) diquarks. The mass of the $cq$ diquark or $\bar{c}\bar{q}$ antidiquark with spin-00 is predicted with ∼$\sim 2175$ MeV, and the spin-$1$ one is predicted with ∼$\sim 2220$ MeV. The masses of hidden charmed tetraquarks from $1S$ to $2P$ excitations are systemically calculated in terms of the same potentials. It is found that the mass of hidden charmed tetraquark without radial excitation grows higher in $1^{+-},~{}1^{++},~{}1^{--},~{}0^{-+},~{}0^{--},~{}1^{-+},~{}\cdots$ sequence, and the tetraquarks with exotic $J^{PC}=0^{--},~{}1^{-+}$ have higher masses. The hidden charmed tetraquarks with radial excitations have masses larger than $4300$ MeV. The $1S-1P$ and $1S-2S$ mass splittings of the hidden charmed tetraquarks are about $390-400$ MeV and $550-570$ MeV, respectively, which are about $70$ MeV and $50$ MeV smaller than those of normal charmonium. The $1P-2P$ and $2S-2P$ mass splittings are similar to those for conventional $c\bar{c}$ charmonium mesons. Based on our predicted masses for hidden charmed tetraquarks, some XYZ exotics are analyzed and tentatively assigned. $X^{*}(3860)$ is possibly the $0^{++}$ tetraquark. $Z_{c}(3900)$ and $X(3940)$ are possibly the $1^{+-}$ tetraquarks, and $X(3872)$ is possibly a $1^{++}$ tetraquark. $X(4250)$ may be a $0^{-+}$, $0^{++}$ or $1^{-+}$ tetraquark, $X(4240)$ may be a $0^{--}$ tetraquark. With radial excitations, $X(4350)$ may be a $0^{++}$ tetraquark, $Z_{c}(4430)$ may be a $1^{+-}$ tetraquark, $X(4630)$ may be a $0^{-+}$ or $1^{-+}$ tetraquark, and $X(Y)(4660)$ may be the $1^{--}$ tetraquark. $Y(4008)$ or $Y(4390)$ seems impossibly the $1^{--}$ tetraquark. In experiments, it is also important to take into account the mixing of normal charmoniums with hidden-flavor charmonium-like exotics for the identification of a normal charmonium or a charmonium-like exotic.
@@ I Introduction
A four-quark state consists of two quarks and two antiquarks. The concept of four-quark state was first mentioned by Gell-Mann, and four-quark states were subsequently studied in hadron scattering amplitudes. The discovery of $X(3872)$ stirred people’s great interest on four-quark states. $X(3872)$ was first observed in exclusive decays $B^{\pm}$ with $X(3872)\to\pi^{+}\pi^{-}J/\Psi$ by Belle [1], where $\pi^{+}\pi^{-}$ is found to be dominated by a $\rho^{0}$ with isospin symmetry violating. Later, the isospin violating channel $X(3872)\to\rho^{0}J/\psi$ [2] was observed. A surprising $\pi^{+}\pi^{-}\pi^{0}J/\Psi$ mode was also reported by Belle [3], but subsequent experiments by Belle [4, 5] did not observe such a $3\pi J/\psi$ decay mode in the radiative decays of $\Upsilon(1S)$ and $\Upsilon(2S)$. However, the isospin conserving channel $X(3872)\to\omega J/\psi$ was observed [3, 6, 7]. The isospin, G parity and $J^{PC}$ quantum numbers of $X(3872)$ have finally been determined with $I^{G}(J^{PC})=0^{+}(1^{++})$ [8].
For its exotic properties different from a conventional $c\bar{c}$ meson, $X(3872)$ has been extensively studied and assigned in all possible models. Many experiments have also been performed for its different features in different facilities [9, 10, 11, 12, 13, 14, 15, 16, 17, 18]. So far, there are many interpretations of $X(3872)$ based on different concerns from different perspectives. $X(3872)$ was assigned with a $(c\bar{q})(q\bar{c})$-like hadronic molecule in Refs. [19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29], and with a tetraquark state consisting of a $cq$ diquark and an $\bar{c}\bar{q}$ antidiquark in Refs. [30, 31, 32, 33, 34, 35, 36, 37, 38]. It was interpreted as a hybrid charmonium in Ref. [39], and as an adjoint charmonium in Ref. [40]. $X(3872)$ was interpreted as a normal $\chi_{c1}(2P)$ charmonium in Ref. [41], and as a mixing state, for example, a normal $c\bar{c}$ charmonium mixing with $D\bar{D}^{*}+\bar{D}D^{*}$ [42] or a diquark-antidiquark mixing with a molecule [43]. It was also interpreted as kinematic or threshold effects from virtual particles [44, 45, 46]. Among these different interpretations, the hadronic molecule interpretation of $X(3872)$ was once most popularly advocated.
There were also some investigations on how to distinguish different interpretations. With a study of the line shape of $X(3872)$, the hadronic molecule interpretation may be distinguished from the virtual charm mesons interpretation [47]. Through the mass spectrum, different patterns for the XYZ states in the hadro-charmonium picture, the tetraquark picture as well as the hadronic molecular picture have been explored [48]. Through an effective range defined through the low energy expansion of the scattering phase, the compact tetraquark nature and the loosely bound molecular one to a state can be discriminated [49]. Polosa et al. suggested that a measure of the ratio $R=\frac{B(X\to\psi^{\prime}\gamma)}{B(X\to\psi\gamma)}$ for radiative decays of $X(3872)$ can serve as a strong indication to discriminate between the molecular and compact interpretations [38]. In $pp$ collisions, some observables were proposed to serve as criterions to identify the $X(3872)$ as either a molecular state or a compact tetraquark state [50].
From the $pp$ collision data, the $X(3872)\to\psi(2S)\gamma$ process is observed for the first time by LHCb collaboration. In particular, the ratio of two partial widths is measured with
Based on the proposal and calculation by Polosa et al. [38], the measured ratio does not support the pure molecule interpretation and strongly indicates a sizeable compact charmonium or tetraquark component within the $X(3872)$.
@@
In addition to $X(3872)$, many hidden charmed exotics have been observed by different experimental collaborations. In Table 1, these hidden charmed exotics (without strange or antistrange quark component) once known as XYZ states in present PDG [8] are listed, where the measured masses and widths are presented. The $J$ and $P$ quantum numbers of $X(3823)$ need confirmation. The new naming scheme for the neutral-flavor hidden charm exotics is employed in the second column. Some of the exotics have been renamed as normal charmoniums in present PDG though these assignments are not definite. The last column indicates the experimental collaborations to the first relevant observations.
@@
In theoretical point of view, the interpretations of exotic states are either based on phenomenological investigations or on effective quantum field theory. In fact, the interpretation of $X(3872)$ or other exotic states depends on the interactions and models employed. How to identify $X(3872)$ among different interpretations is always a puzzle and problem waiting to be solved. So far, there is not a single theoretical model to give a unified understanding of the spectroscopy, decay patterns, production features, and structure of those exotic states. However, the features of many XYZ are unknown yet, the $J^{PC}$ quantum numbers of some exotics are even not determined.
@@
Diquark was first mentioned by Gell-Mann [52], and subsequently introduced to describe a baryon as a composite state of two bodies, a quark and a diquark [53, 54, 55, 56, 57, 58, 59, 60, 61]. In this way, the puzzle of supernumerary baryons could be understood and some features of baryon spectroscopy are well described. Later, the diquark as a fundamental composite was widely employed to construct many kinds of multi-quark systems [62, 63, 64, 65, 66, 30, 67, 68, 69, 70, 71, 72, 73]. The diquark cluster in hadrons is in fact a kind of strong correlation between a pair of quarks. The two quarks attract each other and make up a diquark with a totally antisymmetric symmetry, where the two quarks are also in an antisymmetric color representation. The diquark is either thought as stuff without internal structure or thought as a cluster with a finite size.
@@
Phenomenological evidences for diquarks are mainly on baryon spectroscopy, the $\Delta I={1\over 2}$ rule in weak non-leptonic decays, and some features of parton distribution functions etc.[64, 67]. Recently, positive measure of the production cross sections of hyperons and charmed baryons from $e^{+}e^{-}$ annihilation was performed by Belle collaboration [74]. However, there is not yet any conclusive experimental evidence for the existence of diquark cluster in baryons [75, 76, 77], it is not clear either which system two quarks will make a diquark in.
@@
From a phenomenological analysis [67], the mass difference between good and bad light diquarks is estimated with ∼$\sim 200$ MeV (${2\over 3}$ of the $\Delta$-nucleon mass difference) in light baryons [67]. The mass difference is about $212$ MeV in baryons with one heavy quark where the mass difference between a good light diquark and a light quark is ∼$\sim 310$ MeV. Masses of diquarks have also been calculated in many different methods such as Bethe-Salpeter equation [78, 79, 80, 81], constituent quark model [82, 30, 83, 59, 58, 84, 61, 85], Faddeev equations [86], instanton model [87], QCD sum rules [88, 34, 89, 90], and lattice theory [91, 92].
@@
As well known, quantum chromodynamics (QCD) is the right theory for strong interaction, and every system confined through strong interaction in principle could be understood from QCD or lattice QCD. However, for the limited possibility of computer at present, many topics related to non-perturbation and confinement has to rely on effective theories or phenomenological theories. Constituent quark potential models have made great progresses in the understanding of conventional mesons and baryons, and have been employed to study exotic hadrons.
@@
In constituent quark potential models, the spectra of different tetraquarks have been calculated in a diquark-antidiquark picture in many studies, and every more possible calculation of hidden charmed tetraquarks is expected to. The spectrum of $c\bar{q}c\bar{q}$ tetraquarks was calculated by means of an algebraic mass formula where the mass of $X(3872)$ was set as an input [30]. The tetraquark spectrum was calculated in a relativistic diquark-antidiquark model with one-gluon exchange and long-range vector and scalar linear confining potentials, where the diquarks are considered to have an internal structure with form factor [93]. The spectra of hidden charmed $[qc][\bar{q}\bar{c}]$ and $[sc][\bar{s}\bar{c}]$ tetraquarks were calculated in a relativized diquark model where the diquark masses are extracted from previous studies [37].
@@
In this paper, a hidden charmed tetraquark without strange or antistrange quark component is thought as a composite with a compact $cq$ (q=u, d) diquark and a compact $\bar{c}\bar{q}$ antidiquark. The interactions between a quark and a quark, an antiquark and an antiquark, a quark and an antiquark, and a diquark and an antidiquark are described by universal Semay-Silvestre-Brac potentials [94, 95]. These potentials have been recently employed to calculate the masses of the doubly charmed tetraquarks [85]. In the practical calculation process, the masses of relevant diquarks are first computed in terms of the potentials, and the masses of hidden charmed tetraquarks from $1S$ to $2P$ excitations are subsequently computed. In this way, some features of the mass spectra is obtained for hidden charmed tetraquarks with different inner structure. Based on our numerical results, all the observed exotics in Table 1 are examined, and possible assignments of some XYZ are made.
@@
The work is organized as follows, a simple description of the constituent quark potential model is presented in Sec. II after the introduction. Sec. III is devoted to the practical calculation and analyses of the observed XYZ exotics. The last section gives a brief summary and discussion.
@@ II Constituent quark potential model
For a practical calculation, a modified potential for the quark-antiquark interaction in a norelativistic potential model proposed by Semay and Silvestre-Brac  [94, 95] was employed
@@
$=$ −$-\frac{\alpha(1-e^{-r/r_{c}})}{r}+\lambda r^{p}+C$
@@
The detail of other parameters can be found in Refs. [94, 95]. The quark-quark interaction potential in baryon or multiquark states is similar with $V_{qq}={\frac{1}{2}}V_{q\bar{q}}$.
@@
×$\times\left(\frac{\vec{s}_{d}}{m_{d}}+\frac{\vec{s}_{\bar{d}}}{m_{\bar{d}}}\right)\cdot\vec{L}-\frac{1}{2r}\frac{\partial V_{conf}}{\partial r}\left(\frac{\vec{s}_{d}}{m_{d}^{2}}+\frac{\vec{s}_{\bar{d}}}{m_{\bar{d}}^{2}}\right)\cdot\vec{L}$
@@
×$\times\left(\frac{3\vec{s}_{d}\cdot\vec{r}\vec{s}_{\bar{d}}\cdot\vec{r}}{r^{2}}-\vec{s}_{d}\cdot\vec{s}_{\bar{d}}\right).$
For the calculation of masses of $cq$ diquark and $\bar{c}\bar{q}$ antidiquark, the Hamiltonian is the same as Eq. (2) except that the following quark-quark potential inside the diquark/antidiquark was employed
@@
is composed of color, flavor, spin and spatial wave functions $\chi_{c}$, $\chi_{f}$, $\chi_{s}$ and $\Psi_{lm}(\vec{r})$. Due to the interaction in diquarks, only color $\mathbf{\bar{3}}_{c}(\mathbf{3}_{c})$ diquark (antidiquark) is considered in this work. Therefore two kinds of diquarks/antidiquarks satisfy the symmetry constraints: the flavor symmetric one with spin-$1$ denoted as $\{cq\}^{\bar{3}}_{1}$/$\{\bar{c}\bar{q}\}^{3}_{1}$ and the flavor antisymmetric one with spin-00 denoted as $[cq]^{\bar{3}}_{0}$/$[\bar{c}\bar{q}]^{3}_{0}$. In this way, the constructed wave functions and corresponding quantum numbers of some ground tetraquarks are listed [30, 97, 37]
@@
where $P=(-1)^{L}$ with $L$ indicating the orbital angular momentum between the diquark and the antidiquark. For neutral-flavor hidden charmed tetraquarks, they may have definite $C$ quantum numbers. In a convenient basis for the $C$ parity, the $C$ parity of a tetraquark is defined in Ref. [37] after recoupling of the quark spins
$C$ =(−$=(-1)^{L+s_{c\bar{c}}+s_{q\bar{q}}},$ (7)
@@
For normal neutral-flavor quark-antiquark mesons, possible $J^{PC}$ quantum numbers are $P=(-1)^{L+1}$ and $C=(-1)^{L+S}$. Accordingly, quarkonia has natural $J^{PC}$ quantum numbers such as $0^{-+}$, $0^{++}$, $1^{--}$, $1^{+-}$, $1^{++}$, $2^{--}$, $2^{-+}$, $2^{++}$, ⋯$\cdots$, and cannot have unnatural $J^{PC}$ quantum numbers such as $0^{--}$ and $1^{-+}$ etc. For neutral-flavor tetraquarks, they may have such exotic quantum numbers unnatural for quarkonia.
In order to solve the Schrödinger equation, the spatial wave function $\psi_{lm}(\vec{r})$ is expanded in terms of Gaussian functions as the trial wave function as did in Ref. [98]. Therefore, the Schrödinger equation is transformed into a matrix eigenvalue problem which can be solved as the procedure in Refs. [98, 85].

### v2

@@ Abstract
In a constituent quark model, a hidden charmed tetraquark is assumed consisting of a $cq$ diquark and an $\bar{c}\bar{q}$ antidiquark or vice versa. The Semay-Silvestre-Brac potentials are employed to calculate the masses of $cq$ (q=u, d) diquarks. The mass of the $cq$ diquark or $\bar{c}\bar{q}$ antidiquark with spin-0 is predicted with $\sim 2175$ MeV, and the spin-$1$ one is predicted with $\sim 2220$ MeV. The masses of hidden charmed tetraquarks from $1S$ to $2P$ excitations are systematically calculated in terms of the same potentials. It is found that the mass of hidden charmed tetraquarks without radial excitation grows higher in $1^{+-},~1^{++},~1^{--},~0^{-+},~0^{--},~1^{-+},$$\cdots$ sequence, and the tetraquarks with exotic $J^{PC}=0^{--},~1^{-+}$ have higher masses. The hidden charmed tetraquarks with radial excitations have masses greater than $4300$ MeV. The 1S1P1S-1P and 1S2S1S-2S mass splittings of the hidden charmed tetraquarks are about 390400390-400 MeV and 550570550-570 MeV, respectively, which are about $50$ and $40$ MeV smaller than those of normal charmoniums. The 1P2P1P-2P and 2S2P2S-2P mass splittings are similar to those for conventional $c\bar{c}$ charmonium mesons. Based on our predicted masses for hidden charmed tetraquarks, some XYZ exotics are analyzed and tentatively assigned. $X^{*}(3860)$ is possibly the $0^{++}$ tetraquark. $Z_{c}(3900)$ and $X(3940)$ are possibly the $1^{+-}$ tetraquarks, and $X(3872)$ is possibly a $1^{++}$ tetraquark. $X(4250)$ may be a $^{-+}, 0++0^{++}$ or $1^{-+}$ tetraquark, $X(4240)$ may be a $0^{--}$ tetraquark. With radial excitations, $X(4350)$ may be a $0^{++}$ tetraquark, $Z_{c}(4430)$ may be a $1^{+-}$ tetraquark, $X(4630)$ may be a $0^{-+}$ or $1^{-+}$ tetraquark, and $X(Y)(4660)$ may be the $1^{--}$ tetraquark. $Y(4008)$ or $Y(4390)$ seems impossible to be the $1^{--}$ tetraquark. In experiments, for the identification of a normal charmonium or a charmonium-like exotic, it is also important to take into account the mixing of normal charmoniums with hidden-flavor charmonium-like exotics.
@@ I Introduction
A four-quark state consists of two quarks and two antiquarks. The concept of four-quark states was first mentioned by Gell-Mann, and four-quark states were subsequently studied in hadron scattering amplitudes. The discovery of $X(3872)$ stirred great interest in four-quark states. $X(3872)$ was first observed in exclusive decays $B^{\pm}$ with $X(3872)\to\pi^{+}\pi^{-}J/\Psi$ by Belle [1], where $\pi^{+}\pi^{-}$ is found to be dominated by $\rho^{0}$ with isospin symmetry violating. Later, the isospin violating channel $X(3872)\to\rho^{0}J/\psi$ [2] was observed. A surprising $\pi^{+}\pi^{-}\pi^{0}J/\Psi$ mode was also reported by Belle [3], but subsequent experiments by Belle [4, 5] did not observe such a $3\pi J/\psi$ decay mode in the radiative decays of $\Upsilon(1S)$ and $\Upsilon(2S)$. However, the isospin conserving channel $X(3872)\to\omega J/\psi$ was observed [3, 6, 7]. The isospin, G parity and $J^{PC}$ quantum numbers of $X(3872)$ have finally been determined with $I^{G}(J^{PC})=0^{+}(1^{++})$ [8].
@@
In addition to $X(3872)$, many hidden charmed tetraquark states have been observed by different experimental collaborations. Hidden charmed states are states that contain both a $c$ and a $\bar{c}$ component. In Table 1, the hidden charmed exotics known as XYZ states without strange or antistrange quark component in the present Particle Data Group (PDG) [8] are listed, where the measured masses and widths are presented. The quantum numbers $J$ and $P$ of $X(3823)$ need confirmation. The new naming scheme for the neutral-flavor hidden charm exotics is employed in the second column. Some of the exotics have been renamed as normal charmoniums in the present PDG though these assignments are not definite. The last column indicates the experimental collaborations to the first relevant observations.
@@
For its exotic properties that are different from a conventional $c\bar{c}$ meson, $X(3872)$ has been extensively studied and assigned in all possible models. Many experiments have also been performed for its different features in different facilities [10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. So far, there are many interpretations of $X(3872)$ based on different concerns from different perspectives. $X(3872)$ was assigned with a $(c\bar{q})(q\bar{c})$-like hadronic molecule in Refs. [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30], and with a tetraquark state consisting of a $cq$ diquark and an $\bar{c}\bar{q}$ antidiquark in Refs. [31, 32, 33, 34, 35, 36, 37, 38, 39]. It was interpreted as a hybrid charmonium in Ref. [40], and as an adjoint charmonium in Ref. [41]. $X(3872)$ was interpreted as a normal $\chi_{c1}(2P)$ charmonium in Ref. [42], and as a mixing state, for example, a normal $c\bar{c}$ charmonium mixing with $D\bar{D}^{*}+\bar{D}D^{*}$ [43] or diquark-antidiquark mixing with a molecule [44]. It was also interpreted as kinematic or threshold effects from virtual particles [45, 46, 47]. Among these different interpretations, the hadronic molecule interpretation of $X(3872)$ was the one most popularly advocated.
@@
A diquark was first mentioned by Gell-Mann [48], and subsequently introduced to describe a baryon as a composite state of two bodies, a quark and a diquark [49, 50, 51, 52, 53, 54, 55, 56, 57]. In this way, the puzzle of supernumerary baryons could be understood and some features of baryon spectroscopy were well described. Later, the diquark as a fundamental composite was widely employed to construct many kinds of multiquark systems [58, 59, 60, 61, 62, 31, 63, 64, 65, 66, 67, 68, 69]. The diquark cluster in hadrons is in fact a kind of strong correlation between a pair of quarks. The two quarks attract each other and make up a totally antisymmetric diquark, where the two quarks are also in an antisymmetric color representation. There are two kinds of antisymmetric light diquarks, scalar diquark and vector diquark. The scalar diquark with spin-0 has a lower mass in comparison to the vector diquark with spin-$1$. They are usually called the “good” and “bad” diquarks, respectively, as suggested in Ref. [63]. The diquark is thought either as stuff without internal structure or as a cluster with a finite size.
@@
Phenomenological evidence for diquarks is mainly found in baryon spectroscopy, the $\Delta I={1\over 2}$ rule in weak nonleptonic decays, and some features of parton distribution functions [60, 63]. Recently, a positive measure of the production cross sections of hyperons and charmed baryons from $e^{+}e^{-}$ annihilation was performed by the Belle collaboration [70]. However, there is not yet any conclusive experimental evidence for the existence of diquark cluster in baryons [71, 72, 73], and it is not clear either which system is the right one for two quarks to make a diquark.
@@
From a phenomenological analysis [63], the mass difference between good and bad light diquarks is estimated with $\sim 200$ MeV ($2\over 3$} of the $\Delta$-nucleon mass difference) in light baryons [63]. The mass difference is about $212$ MeV in baryons with one heavy quark where the mass difference between a good light diquark and a light quark is $\sim 310$ MeV. Masses of diquarks have also been calculated in many different methods such as the Bethe-Salpeter equation [74, 75, 76, 77], constituent quark model [78, 31, 79, 55, 54, 80, 57, 81], Faddeev equations [82], instanton model [83], QCD sum rules [84, 35, 85, 86], and lattice theory [87, 88].
There were also some investigations on how to distinguish different interpretations. With a study of the line shape of $X(3872)$, the hadronic molecule interpretation may be distinguished from the virtual charm mesons interpretation [89]. Through the mass spectrum, different patterns for the XYZ states in the hadro-charmonium picture, the tetraquark picture as well as the hadronic molecular picture have been explored [90]. Through an effective range defined through the low energy expansion of the scattering phase, the compact tetraquark nature and the loosely bound molecular one to a state can be discriminated [91]. In the Born-Oppenheimer effective theory, the authors in Ref. [92] claimed that XYZ exotics of any composition with two heavy quarks such as hybrids, tetraquarks, pentaquarks, doubly heavy baryons, and quarkonia can all be addressed. Polosa et al. suggested that a measure of the ratio $R=\frac{B(X\to\psi^{\prime}\gamma)}{B(X\to\psi\gamma)}$ for radiative decays of $X(3872)$ can serve as a strong indication to discriminate between the molecular and compact interpretations [39]. In their calculation, $R\sim 1$ or larger was found for a compact diquark-antidiquark state and $R$ was of order $10^{-2}$ for a molecule state. In $pp$ collisions, some observables were proposed to serve as criteria to identify the $X(3872)$ as either a molecular state or a compact tetraquark state [93].
From the $pp$ collision data, the $X(3872)\to\psi(2S)\gamma$ process is observed for the first time by the LHCb collaboration. In particular, the ratio of two partial widths is measured with
Based on the proposal and calculation by Polosa et al. [39], the measured ratio does not support the pure molecule interpretation and strongly indicates a sizeable compact charmonium or tetraquark component within $X(3872)$.
In the theoretical point of view, the interpretations of exotic states are based on either phenomenological investigations or on effective quantum field theory. In fact, the interpretation of $X(3872)$ or other exotic states depends on the interactions and models employed. How to identify $X(3872)$ among different interpretations is always a puzzle and problem waiting to be solved. So far, there is not a single theoretical model to give a unified understanding of the spectroscopy, decay patterns, production features, and structure of those exotic states. However, the features of many XYZ exotics are still unknown, and the quantum numbers $J$, $P$ and $C$ (if they exist) of some exotics are not determined.
@@
According to current wisdom, quantum chromodynamics (QCD) is the proper theory to describe strong interaction, and every system confined through strong interaction in principle could be understood from QCD or lattice QCD. However, with the limitations of computers at present, many topics related to nonperturbation and confinement have to rely on effective theories or phenomenological theories. Constituent quark potential models have made great progress in the understanding of conventional mesons and baryons, and have been employed to study exotic hadrons.
@@
In constituent quark potential models, the spectra of different tetraquarks have been calculated in a diquark-antidiquark picture in many studies, and more possible calculations of hidden charmed tetraquarks are expected. The spectrum of $c\bar{q}c\bar{q}$ tetraquarks was calculated by means of an algebraic mass formula where the mass of $X(3872)$ was set as an input [31]. The tetraquark spectrum was calculated in a relativistic diquark-antidiquark model with one-gluon exchange and long-range vector and scalar linear confining potentials, where the diquarks are considered to have an internal structure with form factor [94]. The spectra of hidden charmed $qc][\bar{q}\bar{c}]$ and $sc][\bar{s}\bar{c}]$ tetraquarks were calculated in a relativized diquark model where the diquark masses are extracted from previous studies [38].
@@
In this paper, a hidden charmed tetraquark without a strange or antistrange quark component is thought as a composite with a compact $cq$ (q=u, d) diquark and a compact $\bar{c}\bar{q}$ antidiquark. The interactions between a quark and a quark, an antiquark and an antiquark, a quark and an antiquark, and a diquark and an antidiquark are described by universal Semay-Silvestre-Brac potentials [95, 96]. These potentials have been recently employed to calculate the masses of the doubly charmed tetraquarks [81]. In the practical calculation process, the masses of relevant diquarks are first computed in terms of the potentials, and the masses of hidden charmed tetraquarks from $1S$ to $2P$ excitations are subsequently computed. In this way, some features of the mass spectra are obtained for hidden charmed tetraquarks with different inner structures. Based on our numerical results, all the observed exotics in Table 1 are examined, and possible assignments of some XYZ exotics are made.
@@
The work is organized as follows, a simple description of the constituent quark potential model is presented in Sec. II after the introduction. Section III is devoted to the practical calculation and analyses of the observed XYZ exotics. The last section gives a brief summary and discussion.
@@ II Constituent quark potential model
For a practical calculation, a modified potential for the quark-antiquark interaction in a norelativistic potential model proposed by Semay and Silvestre-Brac  [95, 96] was employed, with quark-antiquark interaction potential
@@
$=$ $-\frac{\alpha(1-e^{-r/r_{c}})}{r}+\lambda r^{p}+C$
@@
The details on the parameters can be found in Refs. [95, 96] in which the parameters are determined through a systematic investigation on mesons and baryons. The quark-quark interaction potential in baryon or multiquark states is similar to $V_{qq}={\frac{1}{2}}V_{q\bar{q}}$.
@@
$\times\left(\frac{\vec{s}_{d}}{m_{d}}+\frac{\vec{s}_{\bar{d}}}{m_{\bar{d}}}\right)\cdot\vec{L}-\frac{1}{2r}\frac{\partial V_{conf}}{\partial r}\left(\frac{\vec{s}_{d}}{m_{d}^{2}}+\frac{\vec{s}_{\bar{d}}}{m_{\bar{d}}^{2}}\right)\cdot\vec{L}$
@@
$\times\left(\frac{3\vec{s}_{d}\cdot\vec{r}\vec{s}_{\bar{d}}\cdot\vec{r}}{r^{2}}-\vec{s}_{d}\cdot\vec{s}_{\bar{d}}\right).$
For the calculation of masses of the $cq$ diquark and $\bar{c}\bar{q}$ antidiquark, the Hamiltonian is the same as Eq. (2) except that the following quark-quark potential inside the diquark/antidiquark was employed
@@
is composed of color, flavor, spin and spatial wave functions $\chi_{c}$, $\chi_{f}$, $\chi_{s}$ and $\Psi_{lm}(\vec{r})$. Due to the interaction in diquarks, only color $\mathbf{\bar{3}}_{c}(\mathbf{3}_{c})$ diquark (antidiquark) is considered in this work. Therefore two kinds of diquarks/antidiquarks satisfy the symmetry constraints: the flavor symmetric one with spin-$1$ denoted as $\{cq\}^{\bar{3}}_{1}$$\{\bar{c}\bar{q}\}^{3}_{1}$ and the flavor antisymmetric one with spin-0 denoted as $cq]^{\bar{3}}_{0}$$\bar{c}\bar{q}]^{3}_{0}$. In this way, the constructed wave functions and corresponding quantum numbers of some ground state tetraquarks are listed [31, 98, 38]
@@
where $P=(-1)^{L}$ with $L$ indicating the orbital angular momentum between the diquark and the antidiquark. For neutral-flavor hidden charmed tetraquarks, they may have definite $C$ quantum numbers. On a convenient basis for the $C$ parity, the $C$ parity of a tetraquark is defined in Ref. [38] after recoupling of the quark spins
$C$ $=(-1)^{L+s_{c\bar{c}}+s_{q\bar{q}}},$ (7)
@@
For normal neutral-flavor quark-antiquark mesons, possible $J^{PC}$ quantum numbers are $P=(-1)^{L+1}$ and $C=(-1)^{L+S}$. Accordingly, quarkonia has natural $J^{PC}$ quantum numbers such as $^{-+}, 0++0^{++}$, $^{--}, 1+1^{+-}$, $1^{++}$, $^{--}, 2+2^{-+}$, $2^{++}$, $\cdots$, and cannot have unnatural $J^{PC}$ quantum numbers such as $0^{--}$ and $1^{-+}$. For neutral-flavor tetraquarks, they may have such exotic quantum numbers that are unnatural for quarkonia.
In order to solve the Schrödinger equation, the spatial wave function $\psi_{lm}(\vec{r})$ is expanded in terms of Gaussian functions as the trial wave function as done in Ref. [99]. Therefore, the Schrödinger equation is transformed into a matrix eigenvalue problem which can be solved as the procedure in Refs. [99, 81]. The methods in Ref. [81] developed for doubly charmed tetraquarks are now applied to tetraquarks with hidden charm.